OpenStudy (anonymous):

you purchase a 50-meter roll of fencing, at a cost of $14.95 per meter, in order to build a dog run that is 340 cm wide and 1380 cm long. After you cut the four pieces of fencing from the roll, how much fencing is left on the roll?

OpenStudy (anonymous):

first convert the units - you need all units in cm or m

OpenStudy (ganpat):

i believe dog run would be an rectangle.. so the perimeter of rectangle = 2 (l + b) given l = 1380 and b = 340.. So P = 2 (1380 + 340) = 3440 cm.. converting it to meters ... 34.40 m.. so left over = 50 - 34.40 = 15.60 meters..,

OpenStudy (anonymous):

Thank you

OpenStudy (ganpat):

w/c, but understood ??

OpenStudy (anonymous):

yes, i get it clearly now. thank you so very much.